Grade upto college level Physical Chemistry

How many millilitres of 0.5 M H2SO4 are needed to dissolve 0.5 g copper (II) carbonate?

Amit Saxena
12 Years agoGrade upto college level

1 Answer

Navjyot Kalra
12 Years ago
Use molarity equation to find volume of H2SO4 solution.
CuCO3 + H­2SO4 → CuSO4 + H2O + CO2
63.5+12+48.98 98g
= 123.5 g
∴ For 123.5 gms of Cu(II) carbonate 98 g of H2SO4 are required
For 0.5 gms of Cu(II) carbonate weight of H2SO4 reqd.
= 98 *0.5/123.5 g = 0.39676 g H2SO4
Weight of requiredH2SO4 = 0.39676 g
Weight of solute in grams
= Mol. wt. *Molarity *Volume in mL/1000
0.39676 = 98 *0.5 *V/1000
orV = 0.39676 *1000/90 *0.5 ml
volume of H2SO4solution = 8.097 ml